OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example: (2,1) a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

use a=4-b and sub it into the second equation:\[2(4-b)+3b=11\]\[8+b=11\]\[b=3\]and a=4-b so, a=4-3=1. If your point is represented by (a,b) then your solution is\[(1,3)\]